OpenStudy (anonymous):

What is the sum? 7/2+(-3/5)

OpenStudy (anonymous):

2,9

OpenStudy (anonymous):

That's not a choice

OpenStudy (anonymous):

29/10 and -21/10 and 4/3 41/10

OpenStudy (anonymous):

ja es ist 29/10

OpenStudy (hba):

@jadi Giving direct answers is not allowed -_-

OpenStudy (aravindg):

gosht this is 3rd time i am seeing this question today .Is this of connections academy?

OpenStudy (hba):

You should check the CoC- http://openstudy.com/code-of-conduct

OpenStudy (jiteshmeghwal9):

\[\dfrac{7}{2}-\dfrac{3}{5}={7(5)-3(2)\over10}=?\]

OpenStudy (hba):

@jadi Is that clear ?

OpenStudy (anonymous):

no its not

OpenStudy (jiteshmeghwal9):

@hba @AravindG u r doing ur jobs, keep it up ;)

OpenStudy (aravindg):

thanks @jiteshmeghwal9

OpenStudy (hba):

@jadi I am gonna run in some reports and then you will be suspended.

OpenStudy (jiteshmeghwal9):

yw :)

OpenStudy (hba):

I hope that is understood now @jadi ?

OpenStudy (anonymous):

why are you doing this ?

OpenStudy (hba):

Because this is not allowed on OS.

OpenStudy (anonymous):

ok what is not allowed moreß

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{red}{\text{Welcome To OS :)}}}\]

OpenStudy (hba):

Check the CoC.

OpenStudy (jiteshmeghwal9):

@jadi since u r new member u do not know rules so this is a warning here not to ask for answers try to think on the question. more instructions are in this link :- http://openstudy.com/code-of-conduct

OpenStudy (jiteshmeghwal9):

Click on blue link :)

OpenStudy (anonymous):

I didn't know either Im sorry

OpenStudy (anonymous):

thanks for telling me

OpenStudy (anonymous):

:3c how is asking for things hurting you? if some one wants to fail their classes and fail their future, let them